# Recovering Hardy spaces from optimal domains of integration operators

### v1

@@ Abstract.
We study the optimal domains for bounded Volterra integration operators $T_{g}$ between distinct Hardy spaces $H^{p}$ and $H^{q}$ of the unit ball. It is shown that the intersection of the optimal domains is equal to $H^{p}$ if $p>q$, whereas if $p<q$, we show that this intersection is genuinely larger. In the unit disk, this problem was recently solved for $p=q$ by Bellavita, Daskalogiannis, Nikolaidis and Stylogiannis.
@@
Primary 47B35; Secondary 30H05, 46E20.
S. Eskandari was funded by the postdoctoral scholarship JCK22-0052 from the Kempe Foundations.
@@ 1. Introduction
[…]
Clearly $V$ is always a vector space. If $X$ and $Y$ are quasi-Banach spaces, then $V$ is a quasinormed space equipped with the topology coming from $\|g\|_{V}=\|T_{g}\|_{X\to Y}+|g(0)|$. Moreover, if convergence in both $X$ and $Y$ implies uniform convergence on compact sets (that is, $X$ and $Y$ continuously embed into $H(\mathbb{B}_{n})$), and they both contain the constants, then $V$ is also complete. This fact is proven in [21] in the setting of one complex variable when $T_{g}$ acts $X\to X$, with a slightly less restrictive requirements on $X$. The reader should have no difficulties verifying that it also holds for $T_{g}:X\to Y$ under the assumptions above.
@@
where $d\sigma$ is the normalized surface measure on $\mathbb{S}_{n}$. When $p\geq 1$, the norm $\|\cdot\|_{H^{p}}$ makes $H^{p}$ a Banach space. When $0<p<1$, the space $H^{p}$ is a complete quasi-Banach space. Moreover, in that case $d_{p}(f,g)=\|f-g\|_{H^{p}}^{p}$ is a complete translation invariant metric on $H^{p}$, making it an F-space in the sense it is used by Rudin [20] (they are not locally convex). For basic properties of these spaces, we refer the reader to [13, 23, 24].
For $T_{g}:H^{p}\to H^{q}$, the spaces $V$ are completely known. In fact (see and [18], or [4, 5] for earlier proofs in the case $n=1$),
  • $p=q$: $V=BMOA$;
  • $p<q$: $V=\mathcal{B}^{\alpha}$;
  • $p>q$: $V=H^{r}$.
@@
Given $X$, $Y$, and $g\in V$, we may ask which holomorphic functions $f$ are mapped into $Y$ by $T_{g}$. This vector space of holomorphic functions will be called the optimal domain of $T_{g}$ and denoted
@@
This space $[g:Y]$ can be equipped with the quasinorm $\|f\|_{[g:Y]}=\|T_{g}f\|_{Y}$. The study of optimal domains for integration operators was initiated by Curbera and Ricker [10, 11], who looked into this concept for the Cesáro operator. In the recent years, there has been a number of interesting papers on the topic, see [2, 3, 7], or [6] for meromorphic optimal domains. We mention that all the aforementioned papers are written in the setting of the unit disk. We also remark that our notation differs slightly from them. However, in the context of Volterra operators only, this notation should not confuse the reader. We have chosen it because it makes the notation $[V:Y]$ that is given below more convenient.
From the definitions it follows that we always have $X\subseteq[g:Y]$ for $g\in V$, and if, for instance $X=Y=H^{p}$ and $n=1$, the inclusion is known to be proper [7]. However, another interesting question is whether the intersection of all the optimal domains agrees with $X$. That is, whether $[V:Y]=X$, where
$V:Y]=\bigcap_{g\in V}[g:Y].$
We understand this problem in the following way. While $X$ and $Y$ uniquely define $V$, when is it true that $Y$ and $V$ uniquely define $X$? Or, in other words, can $X$ be recovered from the optimal domains of bounded Volterra operators $X\to Y$.
Before going further, let us offer a simple example. Let $X=A(\mathbb{B_{n}})$ be the ball algebra and $Y=\mathcal{B}^{1}$ be the standard Bloch space. It is easy to see that $V=\mathcal{B}^{1}$, and that for every $g\in V$ one has $H^{\infty}\subseteq[g:\mathcal{B}^{1}]$, so clearly $X\subsetneq[V:Y]$ in that case. But one might argue that an example such as this one is tailor-made for the identity to fail. So what happens in the Hardy spaces?
In [7] it is proven that
$BMOA:H^{p}]=H^{p},\quad p\geq 1,$
@@
which means that $H^{p}$ can be recovered from the optimal domains of bounded $T_{g}:H^{p}\to H^{q}$ for $p>q$.
On the other hand, when $p<q$ and $\alpha=1+n/q-n/p$, we prove
$H^{p}\subsetneq[\mathcal{B}^{\alpha},H^{q}].$
This means that $H^{p}$ cannot be recovered from optimal domains of bounded $T_{g}:H^{p}\to H^{q}$ for $p>q$. Other properties of optimal domains of this type were very recently considered in [3] for $n=1$.
The proofs of these theorems are presented in the next section. We will also study the boundedness of $S_{g}$ between two Hardy spaces, as this is key to studying optimal domains using the approach in [7]. We also discuss some further generalizations and open problems.
@@
2. Main results
We will go straight into the proofs of the results mentioned in the Introduction.
@@ 2. Main results
Let $1<q<p<\infty$ and r=pq/(pq)r=pq/(p-q). Then $H^{p}=[H^{r}:H^{q}]$.
@@
We follow the reasoning in [7]. Let $g\in V(H^{p},H^{q})=H^{r}$, and $f\in[H^{r}:H^{q}]$. Now, $T_{g}f$ belongs to $H^{q}$, or equivalently, $S_{f}g\in H^{q}$. This means by the closed graph theorem (for F-spaces if necessary; see [20]) that $S_{f}:H^{r}\to H^{q}$ is bounded. The claim now follows from the next theorem. ∎
@@
Let $1<q<p<\infty$, r=pq/(pq)r=pq/(p-q) and $g\in H(\mathbb{B}_{n})$. Then $S_{g}:H^{p}\to H^{q}$ is bounded if and only if $g\in H^{r}$. Moreover, if $g(0)=0$, we have
$\|T_{g}\|_{H^{p}\to H^{q}}\asymp\|g\|_{H^{r}}.$
@@
We first consider the sufficiency. In fact, it follows easily from corresponding results for $T_{g}$ and the multiplication operator $M_{g}$. We provide a direct proof, similar to one for $T_{g}$ in [18], by methods that will be used for the necessity. For $\gamma>1$ and $|\zeta|=1$, we denote by $\Gamma(\zeta)=\Gamma_{\gamma}(\zeta)$ the approach region
@@
For the necessity, we proceed with the argument from [17], based on the Kahane and Khinchine inequalities [16], atomic decomposition [18], area function description of Hardy spaces [1, 18], and factorization of tent spaces [8, 9, 17]. The proof follows the outline of either Theorem 7 or Theorem 8 of [17], except that it is simpler. The proof is sketchy in the sense that we do not discuss the preliminaries on tent spaces is this paper; an interested reader can take a look at [17].
@@
We proceed by using subharmonicity and $|1-\langle z,a_{k}\rangle|=1-|a_{k}|^{2}$ for every $z\in D(a_{k},r)$, arriving at the estimate
(2.1) $\int_{\mathbb{S}_{n}}\left(\sum_{{a_{k}}\in\Gamma(\zeta)}|\lambda_{k}|^{2}|a_{k}|^{2}|g(a_{k})|^{2}\right)^{q/2}\,d\sigma(\zeta)\lesssim\|S_{g}\|^{q}\,\|\lambda\|^{q}_{T^{p}_{2}}.$
@@
The first part is the norm $\|\tau\|_{T^{(qs)^{\prime}}_{(2s)^{\prime}}(Z)}.$ By using 2.1 for the second part, to get the upper bound $\|S_{g}\|^{1/s}\|\lambda\|^{1/s}_{T^{p}_{2}}=\|S_{g}\|^{1/s}\|\lambda\|_{T^{ps}_{2s}}$. Altogether we have
@@
By factorization of tent spaces (essentially due to Cohn and Verbitsky [8], but the present variant is taken from [17]) and considering $\mu_{k}=\tau_{k}\lambda_{k}^{1/s}$ as an arbitrary positive element of the space $T^{(rs)^{\prime}}_{1}$. By duality we have $\{\nu_{k}^{1/s}\}\in T^{rs}_{\infty}$. By the discussion and calculation above, this leads to $\|\nu\|_{T^{r}_{\infty}}\lesssim\|S_{g}\|,$ which is what we wanted to prove.
@@
To present the next theorem, we will need to introduce the standard weighted Bergman spaces. Let $\beta>-1$ and $0<p<\infty$. The weighted Bergman space $A^{p}_{\beta}$ consists of $f\in H(\mathbb{B}_{n})$ with
@@
Theorem 3.
Let $1<p<q<\infty$ and $\alpha=1+n/q-n/p$. Then one has $H^{p}\subsetneq[\mathcal{B}^{\alpha}:H^{q}]$.

### v2

@@ Abstract.
We study the optimal domains for bounded Volterra integration operators $T_{g}$ between Hardy spaces $H^{p}$ and $H^{q}$ of the unit ball. It is shown that the optimal domain of a bounded $T_{g}:H^{p}\to H^{q}$ always strictly contains $H^{p}$. Moreover, the intersection of the optimal domains is equal to $H^{p}$ if $p\geq q$, whereas if $p<q$, we show that this intersection is a genuinely larger tent space of holomorphic functions. In the unit disk, this problem was recently solved for $p=q$ by Bellavita, Daskalogiannis, Nikolaidis and Stylogiannis.
@@
Primary 30H10; Secondary 47G10.
S. Eskandari was funded by the postdoctoral scholarship JCK22-0052 from the Kempe Foundations. We would like to thank the anonymous referee for the valuable comments that have improved the paper.
@@ 1. Introduction
$V_{X,Y}=\{g\in H(\mathbb{B}_{n}):T_{g}\text{ is continuous }X\to Y\}.$
Clearly $V_{X,Y}$ is always a vector space. If $X$ and $Y$ are quasi-Banach spaces, then $V_{X,Y}$ is a quasinormed space equipped with the topology coming from $\|g\|_{V_{X,Y}}=\|T_{g}\|_{X\to Y}+|g(0)|$. Moreover, if convergence in both $X$ and $Y$ implies uniform convergence on compact sets (that is, $X$ and $Y$ continuously embed into $H(\mathbb{B}_{n})$), and they both contain the constants, then $V_{X,Y}$ is also complete. This fact is proven in [30] in the setting of one complex variable when $T_{g}$ acts $X\to X$, with a slightly less restrictive requirements on $X$. The reader should have no difficulties verifying that it also holds for $T_{g}:X\to Y$ under the assumptions above.
@@
where $d\sigma$ is the normalized surface measure on $\mathbb{S}_{n}$. When $p\geq 1$, the norm $\|\cdot\|_{H^{p}}$ makes $H^{p}$ a Banach space. When $0<p<1$, the space $H^{p}$ is a complete quasi-Banach space. Moreover, in that case $d_{p}(f,g)=\|f-g\|_{H^{p}}^{p}$ is a complete translation invariant metric on $H^{p}$, making it an F-space in the sense it is used by Rudin [29] (F-spaces need not be locally convex). For basic properties of these spaces, we refer the reader to [16, 33, 34].
For $T_{g}:H^{p}\to H^{q}$, we write $V_{p,q}=V_{H^{p},H^{q}}$. These spaces are completely known. In fact (see and [25], or [5, 6] for earlier proofs in the case $n=1$),
  • $p=q$: $V_{p,q}=BMOA$;
  • $p<q$: $V_{p,q}=\mathcal{B}^{\alpha}$;
  • $p>q$: $V_{p,q}=H^{r}$.
@@
Given $X$, $Y$, and $g\in V_{X,Y}$, we may ask which holomorphic functions $f$ are mapped into $Y$ by $T_{g}$. This vector space of holomorphic functions will be called the optimal domain of $T_{g}$ and denoted
@@
This space $[g:Y]$ can be equipped with the quasinorm $\|f\|_{[g:Y]}=\|T_{g}f\|_{Y}$. The study of optimal domains for integration operators was initiated by Curbera and Ricker [13, 14], who looked into this concept for the Cesáro operator. In the recent years, there has been a number of interesting papers on the topic, see [2, 3, 9], or [8, 20] for meromorphic optimal domains in the context of Hardy and Bergman-Morrey spaces. We mention that all the aforementioned papers are written in the setting of the unit disk. We also remark that our notation differs slightly from them. However, in the context of Volterra operators only, this notation should not confuse the reader. We have chosen it because it makes the notation $V_{X,Y}:Y]$ that is given below more convenient.
From the definitions it follows that we always have $X\subseteq[g:Y]$ for $g\in V_{X,Y}$, and if, for instance $X=Y=H^{p}$ and $n=1$, the inclusion is known to be proper [9]. However, another interesting question is whether the intersection of all the optimal domains agrees with $X$. That is, whether $V_{X,Y}:Y]=X$, where
$V_{X,Y}:Y]=\bigcap_{g\in V_{X,Y}}[g:Y].$
We understand this problem in the following way. While $X$ and $Y$ uniquely define $V_{X,Y}$, when is it true that $Y$ and $V_{X,Y}$ uniquely define $X$? Or, in other words, can $X$ be recovered from the optimal domains of bounded Volterra operators $X\to Y$.
Before going further, let us offer a simple example. Let $X=A(\mathbb{B_{n}})$ be the ball algebra and $Y=\mathcal{B}^{1}$ be the standard Bloch space. It is easy to see that $V_{X,Y}=\mathcal{B}^{1}$, and that for every $g\in V_{X,Y}$ one has $H^{\infty}\subseteq[g:\mathcal{B}^{1}]$, so clearly $X\subsetneq[V_{X,Y}:Y]$ in that case. But one might argue that an example such as this one is tailor-made for the identity to fail. So what happens in the Hardy spaces?
We will prove that the spaces $g:H^{q}]$ are always complete in their respective topologies, and that if $T_{g}:H^{p}\to H^{q}$ is bounded, then $H^{p}$ is strictly contained in $g:H^{q}]$. In [9] it is proven that
(1.1) $BMOA:H^{p}]=H^{p},\quad p\geq 1,$
@@
which means that $H^{p}$ can be recovered from the optimal domains of bounded $T_{g}:H^{p}\to H^{q}$ for $p>q$. On the other hand, when $p<q$ and $\alpha=1+n/q-n/p$, we prove
$H^{p}\subsetneq[\mathcal{B}^{\alpha}:H^{q}],$
and moreover $\mathcal{B}^{\alpha}:H^{q}]$ has a simple description as a tent space. We also establish the (1.1) in for unit ball.
The results indicate, in particular, that $H^{p}$ cannot be recovered from optimal domains of bounded $T_{g}:H^{p}\to H^{q}$ for $p<q$. Other properties of optimal domains of this type were very recently considered in [3] for $n=1$.
In the next section, we focus on the case $p>q$, but we also prove some basic results on optimal domains such as the fact that $g:H^{q}]$ is always a quasi-Banach space (Theorem 1). For the full characterization of the intersection of optimal domains in the case $p>q$, which is Theorem 3, we will also characterize the boundedness of $S_{g}$ between two Hardy spaces. This is key to studying optimal domains using the approach in [9]. The third section is devoted to the study of the case $p<q$, most notably Theorem 7 where we also characterize the intersection of the optimal domains, providing the necessary tent space definitions.
The fourth section deals with the case $p=q$, where the problem was originally studied. We verify that $H^{p}$ is always strictly contained in the corresponding optimal domain, and record as a consquence of this and the previous results that is always the case for a bounded $T_{g}:H^{p}\to H^{q}$ in Corollary 9. Finally, we show that $H^{p}$ can be recovered for the optimal domains of $T_{g}$ for every $p>0$, see Theorem 10. Perhaps surprisingly, this is the most difficult of the three main proofs, and that is why it is saved for the last. This proof, as well as the other two main proofs, makes use of the theory on tent spaces, together with Kahane and Khinchine inequalities and atomic decompositions.
@@
2. The case $p>q$
We will begin by proving that the optimal domains $g:H^{q}]$ for a non-constant holomorphic $g$ are always quasi-Banach spaces, and Banach spaces, if $q\geq 1$. If $g$ is constant, then the optimal domain is trivially all of $H(\mathbb{B}_{n})$, and the construction gives it the trivial topology, making this case meaningless for further analysis.
@@ 2. The case $p>q$
Let $0<q<\infty$ and $g$ be holomorphic and non-constant. Then the optimal domain $g:H^{q}]$ is a quasi-Banach space. It is further a Banach space if $q\geq 1$.
@@
The map $f\mapsto T_{g}f$ is linear. Therefore it is clear that the topology in $g:H^{q}]$ is given by a norm or a quasi-norm, whenever the topology of $H^{q}$ is. We will investigate the completeness of $g:H^{q}]$. The argument is largely the same as in Theorem 1 of [9].
Let $f\in[g:H^{q}]$. Let $Z_{Rg}=\{z\in\mathbb{B}_{n}:Rg(z)=0\}$. For each point $z\notin Z_{Rg}$, we have
$|f(z)|\leq C_{z}\|f\|_{[g:H^{q}]},$
which follows for using the boundedness of the functional $h\mapsto Rh(z)$ on $H^{q}$. The positive constants $C_{z}$ depend on $1/|Rg(z)|$ and $q$ but vary continuously on $\mathbb{B}_{n}\setminus Z_{Rg}$.
Let now $z\in Z_{Rg}$. This case needs some extra effort, since the zeroes of holomorphic functions of several complex variables need not be isolated inside the domain. We proceed as follows. Since $g$ is not constant, there exists $\zeta\in\mathbb{C}^{n}$ with $|\zeta|=1$, and $\varepsilon>0$, so that the one variable analytic slice $h_{\zeta}(w)=Rg(z+w\zeta)$ satisfies $|h_{\zeta}(w)|>\varepsilon$ when $r<|w|<R$ for some $0<r<R$. We can use subharmonicity of $\log|h_{\zeta}|$ to arrive at the an upper bound for $|f(z)|$ similar to that in [9]. This upper bound can be chosen to be uniform in some neighborhood of $z$ (regardless if the new point belong to $Z_{Rg}$ or not).
For these observations, it follows that Cauchy sequences in $g:H^{q}]$ converge to a holomorphic function uniformly on compact subsets. Now that this has been achieved, the rest of the argument in the cited paper carries over. ∎
Let $0<q<p<\infty$ and r=pq/(pq)r=pq/(p-q). We show that if $T_{g}:H^{p}\to H^{q}$ is bounded, then $H^{p}$ is always strictly contained in the optimal domain $g:H^{q}]$.
Proposition 2.
Let $0<q<p<\infty$, $g\in H^{r}$. Then $H^{p}\subsetneq[g:H^{q}]$.
Proof.
One can argue as in [9]. Indeed, for any $g\in H^{r}$, it is trivially true that $H^{p}$ is contained in $g:H^{q}]$. If there were a $g\in H^{r}$ with $H^{p}=[g:H^{q}]$, the identity map would be bounded and onto $H^{p}\to[g:H^{q}]$. The open mapping theorem (which holds also for F-spaces [29]) then implies that $T_{g}:H^{p}\to H^{q}$ is bounded below.
Since $H^{p}$ is a subspace of $g:H^{q}]$ and embeds there boundedly, if there was a $g\in H^{r}$ with $H^{p}=[g:H^{q}]$, one would have that $T_{g}:H^{p}\to H^{q}$ is bounded below. It is well-known [25] that for this range of $p$ and $q$, $T_{g}:H^{p}\to H^{q}$ is bounded if and only if it is compact. A compact operator cannot be bounded below. ∎
We will now proceed with the proofs of the results mentioned in the Introduction.
Theorem 3.
Let $0<q<p<\infty$ and r=pq/(pq)r=pq/(p-q). Then $H^{p}=[H^{r}:H^{q}]$.
Proof.
We follow the reasoning in [9]. Let $g\in V_{p,q}=H^{r}$, and $f\in[H^{r}:H^{q}]$. Now, $T_{g}f$ belongs to $H^{q}$, or equivalently, $S_{f}g\in H^{q}$. This means by the closed graph theorem (for F-spaces if necessary; see [29]) that $S_{f}:H^{r}\to H^{q}$ is bounded. The claim now follows from the next theorem. ∎
@@
Let $0<q<p<\infty$, r=pq/(pq)r=pq/(p-q) and $g\in H(\mathbb{B}_{n})$. Then $S_{g}:H^{p}\to H^{q}$ is bounded if and only if $g\in H^{r}$. Moreover, if $g(0)=0$, we have
$\|S_{g}\|_{H^{p}\to H^{q}}\asymp\|g\|_{H^{r}}.$
@@
We first consider the sufficiency. In fact, it follows easily from the corresponding results for $T_{g}$ and the multiplication operator $M_{g}$. We provide a direct proof, similar to one for $T_{g}$ in [25], by methods that will be used for the necessity. For $\gamma>1$ and $|\zeta|=1$, we denote by $\Gamma(\zeta)=\Gamma_{\gamma}(\zeta)$ the approach region
@@
For the necessity, we proceed with the argument from [23], based on the Kahane and Khinchine inequalities [21], atomic decomposition [12, 25], area function description of Hardy spaces [1, 25], and factorization of tent spaces [10, 11, 23]. The proof follows the outline of either Theorem 7 or Theorem 8 of [23], except that it is simpler. The proof is sketchy in the sense that we do not discuss the preliminaries on tent spaces is this paper; an interested reader can take a look at [23].
@@
We proceed by using subharmonicity and $|1-\langle z,a_{k}\rangle|=1-|a_{k}|^{2}$ for every $z\in B(a_{k},r)$, arriving at the estimate
(2.1) $\int_{\mathbb{S}_{n}}\left(\sum_{{a_{k}}\in\Gamma(\zeta)}|\lambda_{k}|^{2}|a_{k}|^{4}|g(a_{k})|^{2}\right)^{q/2}\,d\sigma(\zeta)\lesssim\|S_{g}\|^{q}\,\|\lambda\|^{q}_{T^{p}_{2}}.$
@@
The first part is the norm $\|\tau\|_{T^{(qs)^{\prime}}_{(2s)^{\prime}}}.$ By using 2.1 for the second part, to get the upper bound $\|S_{g}\|^{1/s}\|\lambda\|^{1/s}_{T^{p}_{2}}=\|S_{g}\|^{1/s}\|\lambda\|_{T^{ps}_{2s}}$. Altogether we have
@@
By factorization of tent spaces (essentially due to Cohn and Verbitsky [10], but the present variant is taken from [23]) and considering $\mu_{k}=\tau_{k}\lambda_{k}^{1/s}$ as an arbitrary positive element of the space $T^{(rs)^{\prime}}_{1}$, using duality we have $\{\nu_{k}^{1/s}\}\in T^{rs}_{\infty}$. By the discussion and calculation above, this leads to $\|\nu\|_{T^{r}_{\infty}}\lesssim\|S_{g}\|,$ which is what we wanted to prove.
@@
3. The case $p<q$
If $0<p<q<\infty$ and $T_{g}:H^{p}\to H^{q}$ is compact, then Proposition 3.19 of [3] states that $H^{p}\subsetneq[g:H^{q}]$. In fact, the result in the cited paper is formulated on the unit disk, but extends to several complex variables without extra efforts. In this section we will show that the same is true for the intersection of the optimal domains, even when taken over all bounded $T_{g}$. Morever, we completely describe this intersection.
As an appetiser, we will present a short proof that the intersection of the optimal domains is always larger than $H^{p}$, when $p<q$. We will need to introduce the standard weighted Bergman spaces. Let $\beta>-1$ and $0<p<\infty$. The weighted Bergman space $A^{p}_{\beta}$ consists of $f\in H(\mathbb{B}_{n})$ with
@@ 3. The case $p<q$
Proposition 5.
Let $0<p<q<\infty$ and $\alpha=1+n/q-n/p$. Then one has $H^{p}\subsetneq[\mathcal{B}^{\alpha}:H^{q}]$.